Mindy is building a patio that is in the shape of a square. The patio will cover 56 square yards. Find the length of a side of the patio to the nearest tenth of a yard.
Answer by stanbon(75887)

Mindy is building a patio that is in the shape of a square. The patio will cover 56 square yards. Find the length of a side of the patio to the nearest tenth of a yard.
----
Area of a square = (side)^2
------
Solve: 56 = s^2
s = sqrt(56) = 7.4833
==========================
Cheers,
Stan H.
==========